Sol Summary Report – March 2th

Crew 275 Sol Summary Report 02Mar2023

Sol: 18

Summary Title: When space exploration brings philosophy to the table

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With the precipitation of the last few days, we knew that the ground would be too muddy today to conduct EVAs, which is why we decided to wait a bit before performing our second EVA to Candor Chasma. It was apparently the right choice, as we can see from our windows that the mud is quickly drying out.

We conducted routine operations in the Hab, the GreenHab, and the Science Dome, finalizing most of the Human Factors experiments planned for Week 3. With the sunny weather, our Crew Astronomer was finally able to take his first pictures of the Sun. He has been waiting for this moment for so long and is very happy with the result!

The end of the afternoon was dedicated to cleaning the different modules, as we want to make a good impression on the BBC tomorrow. They have also sent us a list of questions that we have started to prepare for tomorrow’s interviews. Some of these questions were actually the topic of a very interesting discussion that we all had together after lunch.

Look Ahead Plan:

Tomorrow (Sol 19), we would like to have our second EVA to Candor Chasma. We then plan to go back there on Saturday (Sol 20). The EVA crews will have to find specific checkpoints based on a 2D map or a 3D render, respectively. We would like to take the opportunity of these EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather. In addition, since no precipitation is expected, we would like to re-install the LOAC next to the other atmospheric instruments.

The BBC is planned to arrive at the MDRS tomorrow (Sol 19) between 8:30 and 9:00. They will follow us on our EVA to Candor Chasma, then film inside the station in the afternoon, and leave at around 17:00.

Anomalies in work: None.

Weather: Sunny, low -1°C/30°F, high 7°C/44°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Request, Operations Report, Journalist Report, GreenHab Report, Astronomy Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site